OpenStudy (anonymous):

Find the equation of a line through (4, -3) that is parallel to the line that passes through these points (-4,7) and (2,4). Write your answer in standard form (Ax + By = C). Show all your work.

OpenStudy (anonymous):

The first step would be to find the slope of a line passing through (-4,7) and (2,4). You can do this with the slope equation \[slope = \frac{ y _{2}-y _{1} }{ x _{2}-x_{1}}\]Plugging in those points would yield (7-4)/(-4-2) which gives -1/2. Lines that are parallel have the same slope so your new line needs the slope -1/2. You can use the formula \[y-y _{1} = m(x-x _{1})\] with y1 = -3 and x1 = 4 and m = -1/2 This will give you y = (-1/2)x+5. In Ax + By = C form, this is x+2y=-2.

OpenStudy (anonymous):

thanks so x+2y=-2 is the answer riggt